How many 7 digit phone numbers are possible if the first digit cannot be 0 or 1 and the first 3 digits cannot be 911 or 411?
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52781)

.

In total, there are 10,000,000  7-digit numbers, starting from 0000000 and completing by 9999999.


Of them, all starting from "0" are prohibited; in all, the amount of these numbers is 1,0000,000.


Next, the numbers starting from "1", are prohibited, too; the amount of these numbers is 1,000,000, again.


The numbers starting from 911 are prohibited; the amount of these numbers is 10,000.


The numbers starting from 411 are prohibited; the amount of these numbers is 10,000.


Therefore, to answer the question, we need subtract 1,000,000, then 1,000,000 again, then 10,000  and then 10,000 again, from 10,000,000.


It will give you the answer


    10,000,000 - 1,000,000 - 1,000,000 - 10,000 - 10,000 = 7,980,000.



--------------
Comment from student: Thank you so much! Your solution is the one that I got. But the answer on the test
answer key is 7*9*9*10*10*10*10. I cannot figure out how that is the correct answer. Any idea?
--------------


My response :

I have two ideas.


    My 1-st idea is that my solution is correct.


    My 2-nd idea is an advise to you: do not accept any other answer.

        Since if the answer is different, it means that the other solution is incorrect (wrong).



Answer by greenestamps(13200)



The answer from the answer key -- 7*9*9*10*10*10*10 -- makes no sense. There is no way to solve the problem simply by finding the numbers of choices for each of the 7 digits.

There are no restrictions on the last four digits; there are 10 choices for each of those.

For the first three digits, all the 3-digit numbers 200 to 999 except for 411 and 911 are allowed. That's 800-2 = 798 possibilities for the first three digits.

So the number of 7 digit phone numbers is 798*10^4 = 7,980,000.

Your answer is correct; the answer key is wrong.